OpenStudy (anonymous):

Write a recursive formula; 4,12,36,108 Formula: an=a n-1 a1=4

OpenStudy (anonymous):

a(n) = 3 a(n-1) a(1) = 4

OpenStudy (ybarrap):

Just a little more detail on how we arrive at this: n=1 a1=4=2^2 n=2, a2=12 = 3*4=3*2^2=3*a1 n=3, a3=36 = 3*12=3*3*4=3^2*2^2=3*a2 n=4, a4=108 = 3*36=3*3*12=3^2*3*4=3^3*2^2=3*a3 Pattern a1=4,an=3*a(n-1) [Note also that any an can be found given an n using, an=3^(n-1) * 4 without knowing the previous value.]